# The jittering jets explosion mechanism (JJEM) in electron capture supernovae

### v1

@@ Abstract
We conduct one-dimensional stellar-evolution simulations of stars with zero age main sequence masses of $M_{\rm ZAMS}=8.8-9.45M_{\odot}$ towards core collapse by electron capture, and find that the convective zone of the pre-collapse core can supply the required stochastic angular momentum fluctuations to set a jet-driven electron capture supernova (ECSN) explosion in the frame of the jittering jets explosion mechanism (JJEM). By our criteria of minimum convective specific angular momentum and an accreted mass during jet-launching of $M_{\rm acc}\simeq 0.001-0.01M_{\odot}$, the layer in the convective zone that when accreted launches the exploding jittering jets resides in the helium-rich zone. This exploding layer is accreted at about a minute to a few hours, depending on the model, after core collapse occurs, much shorter than the time the exploding shock crosses the star. The final (gravitational) mass of the neutron star (NS) remnant is in the range of $M_{\rm NS}=1.25-1.43M_{\odot}$. Our results add to the growing support of the JJEM as the main explosion mechanism of massive stars.
@@ 1 INTRODUCTION
Stars in the intermediate mass range $\Delta M_{\rm ECSN}$ between those that leave a white dwarf remnant and those that undergo FeCCSN, $8~{}M_{\odot}\lesssim M_{\rm{ZAMS}}\lesssim 10~{}M_{\odot}$ where the limits depend on the initial metallicity (e.g., Doherty et al. 2015, 2017; Chanlaridis et al. 2022; Cinquegrana et al. 2023), are presumed to collapse as an electron capture supernova (ECSNe, e.g., Miyaji et al. 1980). The actual mass range for ECSNe is smaller than the above range, i.e., $\Delta M_{\rm ECSN}<1M_{\odot}$. ECSNe do not develop a massive iron core but rather explode when the main isotopes of their core are ${}^{16}$, ${}^{20}$e, ${}^{28}$i, ${}^{32}$, and some ${}^{24}$g; iron group elements account for less than half of the core mass. The electron capture processes on ${}^{24}$g and then ${}^{20}$e disturb the hydrostatic equilibrium in the inner core and the core collapses to form a neutron star in an ECSNe (e.g.,Miyaji et al. 1980; Nomoto 1984, 1987; Poelarends et al. 2008; Leung & Nomoto 2019; Zha et al. 2019; Guo et al. 2024; Limongi et al. 2024). The typical explosion energy of ECSNe is expected to be $E_{\rm ex}\approx 10^{49}-10^{50}{~{}\rm erg}$ and they leave an NS remnant. An example is probably the Crab Nebula with an explosion energy of $E_{\rm ex}\simeq 10^{50}{~{}\rm erg}$ (see analysis by Yang & Chevalier 2015).
@@
In CCSNe (both FeCCSNe and ECSNe) the explosion energy comes from a small fraction of the gravitational energy that the collapsing core releases. There are two alternative mechanisms to channel a small fraction of the gravitational energy to explode the star, the delayed neutrino explosion mechanism and the jittering jets explosion mechanism (JJEM). The collapsing core forms a central proto-NS at nuclear densities that stops the collapse, i.e., the core bounce phase, and drives a shock wave that propagates out. In FeCCSNe the shock stalled at $r\simeq 150{~{}\rm km}$.
According to the delayed neutrino explosion mechanism, heating by neutrinos over a second to several seconds revives the shock, and the re-expanding shock explodes the star (Bethe & Wilson 1985, with hundreds of papers in the four decades since then, e.g., Heger et al. 2003; Janka 2012; Nordhaus et al. 2012; Gabay et al. 2015; Ertl et al. 2016; Müller et al. 2019; Bruenn et al. 2020; Bollig et al. 2021; Burrows et al. 2020; Stockinger et al. 2020a; Fujibayashi et al. 2021; Fryer et al. 2022; Nakamura et al. 2022; Olejak et al. 2022; Boccioli et al. 2023; Burrows et al. 2023).
@@
In the JJEM (Soker, 2010), on the other hand, jets that the newly born neutron star launches in stochastically varying directions, i.e., jittering jets, explode the star (e.g., Papish & Soker 2011, 2014; Gilkis & Soker 2014, 2015, 2016; Quataert et al. 2019; Soker 2019, 2020, 2023a, 2023b; Shishkin & Soker 2021, 2022, 2023; Antoni & Quataert 2022, 2023). Neutrino heating does play a role in the JJEM by boosting the jet-driven explosion (Soker, 2022). The intermittent accretion disks that launch the jittering jets are formed by the accreted gas onto the newly born NS that has (almost) stochastically varying angular momentum, both in magnitude and in direction. The initial source of the stochastic angular momentum is the stochastic motion of convective layers in the pre-collapse core that introduces seed angular momentum perturbations. The angular momentum fluctuations due to the random velocities of the convective cells are amplified by instabilities behind the stalled shock at $r\lesssim 100{~{}\rm km}$ above the newly born NS. The spiral standing accretion shock instability (spiral SASI) is the dominant instability that we expect to amplify the seed angular momentum perturbations of the accreted convective layers (for the SASI see, e.g., Blondin & Mezzacappa 2007; Iwakami et al. 2014; Kuroda et al. 2014; Fernández 2015; Kazeroni et al. 2017; for spiral-SASI see, e.g., Andresen et al. 2019; Walk et al. 2020; Nagakura et al. 2020; Shibagaki et al. 2021).
@@
To form an accretion disk around a NS the specific angular momentum of the accreted matter should be $j_{\rm acc}\gtrsim j_{\rm d,NS}\simeq 2\times 10^{16}{~{}\rm cm}^{2}{~{}\rm s}^{-1}$ (e.g. Papish et al. 2015). Several convective cells contribute the material to the accreted gas at any given time. At a large fraction of the accretion process, according to the JJEM, the specific angular momentum after amplification by the spiral-SASI is sufficiently large to form a short-lived ($\simeq 0.01-0.1{~{}\rm s}$) accretion disk around the NS. We define the angular momentum parameter as
@@
where $r$ is the radius of the layer in the pre-collapse core and $v_{\rm conv}$ is the convective velocity as the one-dimensional stellar evolutionary code gives according to the mixing length theory (MLT). Therefore, the specific angular momentum of the gas formed by several cells will be smaller than $j$, which is the value of one convective cell. The amplification of the specific angular momentum fluctuations by the spiral-SASI ensures that, at sporadic periods, the specific angular momentum of the accreted gas is sufficiently large to form an accretion disk (or an accretion belt) around the NS. Therefore, for launching jittering jets, the value of $j$ can be smaller than $j_{\rm d,NS}$. We follow Shishkin & Soker (2022) and take the constraint to be $j\gtrsim 0.1-0.25j_{\rm d,NS}\simeq 2\times 10^{15}-5\times 10^{15}{~{}\rm cm}^{2}{~{}\rm s}^{-1}$. Earlier studies, in particular, Shishkin & Soker (2022), concluded that progenitors of FeCCSNe have sufficiently large angular momentum fluctuations, i.e., $j>2\times 10^{15}{~{}\rm cm}^{2}{~{}\rm s}^{-1}$, in their convective zones near silicon burning and/or oxygen burning to support the JJEM.
Motivated by recent conclusions that many CCSN remnants are shaped by jets with varying directions, most likely related to the jets that exploded the star in the JJEM (for a review see Soker 2024), we examine whether the pre-collapse convective layers in the outer core and above the core of progenitors of ECSNe can supply sufficiently large angular momentum fluctuations as the JJEM requires. In section 2 we describe our numerical setting and assumptions, and in sections 3 - 5 we describe the results for several stellar models. We summarize this study in section 6
@@ 2 Numerical Scheme
We use a custom large nuclear network of 132 isotopes up to ${}^{60}$n (see Fig. 7 in Appendix) to account for weak reactions, namely, electron capture and Urca processes, that arise after the formation of the ONeMg core (e.g., Jones et al. 2013; Takahashi et al. 2013). The weak rates for the atomic mass of A=20 to A=28 isotopes tabulated by Suzuki et al. (2016) are used to account for reactions relevant to high-density ONeMg cores.
@@ 3.1 General properties
We examine the properties of the convection layer in the outer parts of the core and above the core as close as we can get the simulation to electron-capture core collapse. We do not study the entire chain of events that lead to the collapse of the core which can be found in other studies (e.g., Jones et al. 2013; Woosley & Heger 2015; Takahashi et al. 2013; Nomoto 1987; Miyaji et al. 1980; Leung & Nomoto 2019; Zha et al. 2019; Limongi et al. 2024). These, and other studies, find that in the late evolutionary stages of an ECSN progenitor close to collapse, processes driven by the weak nuclear force, such as the Urca processes and electron capture (mainly onto ${}^{24}$g), remove energy from the center of the core. As a result, the core contracts, and oxygen and neon start to burn off-center, producing heavier elements, mainly ${}^{28}$i and ${}^{32}$. The O+Ne off-center burning can be described as a laminar flame propagating inwards (for treatments of laminar flames see, e.g., Timmes & Swesty 2000; Leung & Nomoto 2019; Jones et al. 2013, 2016). As the deflagration flame approaches the center, it triggers further electron capture reactions (mainly onto ${}^{20}$e) and induces collapse (Takahashi et al. 2013; Woosley & Heger 2015). The burning can reach iron peak elements (e.g., Leung & Nomoto 2019; Takahashi et al. 2013), but ECSNe explode before they develop a massive iron core. Note that during this process, the contraction of the core is in constant competition with oxygen or sometimes neon deflagration, which could lead to a thermonuclear runaway (e.g., Miyaji et al. 1980; Nomoto 1987; Leung & Nomoto 2019).
@@
Due to numerical difficulties, we do not reach actual core collapse. Based on the core composition at the last simulation time, and comparison to Doherty et al. (2015), Jones et al. (2013), Stockinger et al. (2020a), and Limongi et al. (2024) that we referred to above, we infer that the relevant convective zones (see below) above the core in our last stellar model are not likely to change much until the full collapse. There is still some evolution of the inner core to collapse, like towards nuclear statistical equilibrium (NSE) in the inner core. The $8.8~{}M_{\odot}$ model with a solar composition that Stockinger et al. (2020a) present just at core collapse has an inner mass of $0.45~{}M_{\odot}$ in NSE, and a CO core radius of only $\simeq 1200{~{}\rm km}$. Also, they do not have an extended helium-rich layer as we have here, of about $1~{}M_{\odot}$. However, the exact structure depends on the helium core mass. Hillebrandt et al. (1984) find that in their simulation starting with a helium core mass of $2.2~{}M_{\odot}$, at a time of 100 years before the collapse, there is a very thin helium-rich layer, while when they start with a helium core mass of $2.4~{}M_{\odot}$ there is a thick helium layer, $\simeq 1~{}M_{\odot}$ at that time. Our helium core mass is $\simeq 2.4~{}M_{\odot}$, and therefore we expect to have a thick helium layer, as we do get here.

### v2

@@ Abstract
We conduct one-dimensional stellar-evolution simulations of stars with zero-age main sequence masses of $M_{\rm ZAMS}=8.8-9.45M_{\odot}$ towards core collapse by electron capture and find that the convective zone of the pre-collapse core can supply the required stochastic angular momentum fluctuations to set a jet-driven electron capture supernova (ECSN) explosion in the frame of the jittering jets explosion mechanism (JJEM). By our assumed criteria of minimum convective specific angular momentum and an accreted mass during jet-launching of $M_{\rm acc}\simeq 0.001-0.01M_{\odot}$, the layer in the convective zone that when accreted launches the exploding jittering jets resides in the helium-rich zone. Depending on the model, this exploding layer is accreted at about a minute to a few hours after core collapse occurs, much shorter than the time the exploding shock crosses the star. The final (gravitational) mass of the neutron star (NS) remnant is in the range of $M_{\rm NS}=1.25-1.43M_{\odot}$.
@@ 1 INTRODUCTION
Stars in the intermediate mass range $\Delta M_{\rm ECSN}$ between those that leave a white dwarf remnant and those that undergo FeCCSN, $8~{}M_{\odot}\lesssim M_{\rm{ZAMS}}\lesssim 10~{}M_{\odot}$ where the limits depend on the initial metallicity (e.g., Doherty et al. 2015, 2017; Chanlaridis et al. 2022; Cinquegrana et al. 2023), are presumed to collapse as an electron capture supernova (ECSNe, e.g., Miyaji et al. 1980). The actual mass range for ECSNe is smaller than the above range, i.e., $\Delta M_{\rm ECSN}<1M_{\odot}$. ECSNe do not develop a massive iron core but rather explode when the main isotopes of their core are 16O, 20Ne, 28Si, 32S, and some 24Mg; iron group elements account for less than half of the core mass. The electron capture processes on 24Mg and then 20Ne disturb the hydrostatic equilibrium in the inner core and the core collapses to form a neutron star in an ECSNe (e.g., Miyaji et al. 1980; Nomoto 1984, 1987; Poelarends et al. 2008; Leung & Nomoto 2019; Zha et al. 2019; Guo et al. 2024; Limongi et al. 2024). The typical explosion energy of ECSNe is expected to be $E_{\rm ex}\approx 10^{49}-10^{50}{~{}\rm erg}$ and they leave an NS remnant. An example is probably the Crab Nebula with an explosion energy of $E_{\rm ex}\simeq 10^{50}{~{}\rm erg}$ (see analysis by Yang & Chevalier 2015).
@@
In CCSNe (both FeCCSNe and ECSNe), the explosion energy comes from a small fraction of the gravitational energy that the collapsing core releases. The collapsing core forms a central proto-NS at nuclear densities that stops the collapse, i.e., the core bounce phase, and drives a shock wave that propagates out. In FeCCSNe the shock stalled at $r\simeq 150{~{}\rm km}$. There are two alternative mechanisms to channel a small fraction of the gravitational energy to explode the star, the delayed neutrino explosion mechanism and the jittering jets explosion mechanism (JJEM), as we summarize in section 2.
Motivated by recent conclusions (e.g., Bear & Soker 2024; Soker 2024a) that many CCSN remnants are shaped by jets with varying directions, most likely related to the jets that exploded the star in the JJEM (section 2; for a review see Soker 2024b), we examine whether the pre-collapse convective layers in the outer core and above the core of progenitors of ECSNe can supply sufficiently large angular momentum fluctuations as the JJEM requires. In section 3, we describe our numerical setting and assumptions, and in sections 4 - 6, we describe the results for several stellar models. We summarize this study in section 7
2 The two explosion mechanisms
According to the delayed neutrino explosion mechanism, heating by neutrinos over a second to several seconds revives the shock; the re-expanding shock explodes the star (Bethe & Wilson 1985, with hundreds of papers in the four decades since then, e.g., Heger et al. 2003; Janka 2012; Nordhaus et al. 2012; Gabay et al. 2015; Ertl et al. 2016; Müller et al. 2019; Bruenn et al. 2020; Bollig et al. 2021; Burrows et al. 2020; Stockinger et al. 2020; Fujibayashi et al. 2021; Fryer et al. 2022; Nakamura et al. 2022; Olejak et al. 2022; Boccioli et al. 2023; Burrows et al. 2023; Boccioli & Roberti 2024). The neutrino-driven explosion mechanism encounters several challenges. One is that research groups disagree on both qualitative and quantitative results. [^1] In a review talk in the Transients Down Under meeting, held in Melbourne, Australia on 29 January 2024, Hans-Thomas Janka writes in the summary of his talk about the delayed neutrino explosion mechanism: “3D models of different groups disagree: there are qualitative and quantitative differences!”. For instance, Radice et al. (2017) find an explosion time of $t\simeq 0.6{~{}\rm s}$ after the collapse with a leftover proto-neutron star mass of $\rm{M_{NS}}\simeq 1.2~{}M_{\odot}$ and explosion energy to range from $1.2-1.8\times 10^{50}{~{}\rm erg}$ for the models that do explode. For the same ECSN progenitor model of Nomoto (1984, 1987), simulations by other groups have resulted in explosion energies that are substantially different from Radice et al. (2017), e.g., Wanajo et al. (2003) find explosion energies of $0.2-35\times 10^{50}{~{}\rm erg}$, Kitaura et al. (2006) find $\approx 10^{50}{~{}\rm erg}$, and Janka et al. (2008) $<10^{50}{~{}\rm erg}$. Moreover, not all of the models provided by the neutrino-driven manage to explode. For instance, out of the 24 models in Radice et al. (2017), only 15 of them exploded, with the mass range from $\rm{~{}M_{\odot}}=9.0,10.0,11.0$ not exploding in the 1D simulations. Another challenge of the neutrino-driven explosion mechanism is to explain the point-symmetric morphology that was identified in the last year in several CCSNe remnants, and which suggests multi-pair jet activity at explosion (e.g., Bear & Soker 2024; Soker 2024a).
@@ 2 The two explosion mechanisms
In the JJEM (Soker, 2010), on the other hand, jets that the newly born neutron star launches in stochastically varying directions, i.e., jittering jets, explode the star (e.g., Papish & Soker 2011, 2014; Gilkis & Soker 2014, 2015, 2016; Quataert et al. 2019; Soker 2019a, 2020a, 2023, 2024c; Shishkin & Soker 2021, 2022, 2023; Antoni & Quataert 2022, 2023). Neutrino heating does play a role in the JJEM by boosting the jet-driven explosion (Soker, 2022). The accreted gas with stochastically varying angular momentum, in both magnitude and direction, forms the intermittent accretion disks that launch the jittering jets from the newly born NS. The initial source of the stochastic angular momentum is the stochastic motion of convective layers in the pre-collapse core that introduces seed angular momentum perturbations. The angular momentum fluctuations due to the random velocities of the convective cells are amplified by instabilities behind the stalled shock at $r\lesssim 100{~{}\rm km}$ above the newly born NS. The spiral standing accretion shock instability (spiral SASI) is the dominant instability that we expect to amplify the seed angular momentum perturbations of the accreted convective layers (for the SASI see, e.g., Blondin & Mezzacappa 2007; Iwakami et al. 2014; Kuroda et al. 2014; Fernández 2015; Kazeroni et al. 2017; for spiral-SASI see, e.g., Andresen et al. 2019; Walk et al. 2020; Nagakura et al. 2020; Shibagaki et al. 2021). Other types of instabilities are possible. However, to fully explore the instabilities and processes that operate in the JJEM, the simulations must include magnetic fields (e.g., Soker 2018, 2019b, 2020b).
Backed-up by the new finding of multi-pair jet activity in CCSN remnants (e.g., Soker 2024b, a), we here examine whether the criteria that were used for the operation of the JJEM in FeCCSNe also apply to ECSNe.
@@
To form an accretion disk around an NS, the specific angular momentum of the accreted matter $j_{\rm acc}$ should be about equal or larger than the specific angular momentum of a particle in a centrifugally-supported accretion disk on the surface of the NS $j_{\rm d,NS}$, i.e., $j_{\rm acc}\gtrsim j_{\rm d,NS}\simeq 2\times 10^{16}{~{}\rm cm}^{2}{~{}\rm s}^{-1}$ (e.g. Papish et al. 2015). Several convective cells contribute the material to the accreted gas at any time. At a large fraction of the accretion process, according to the JJEM, the specific angular momentum after amplification by the spiral-SASI is sufficiently large to form a short-lived ($\simeq 0.01-0.1{~{}\rm s}$) accretion disk around the NS. We define the angular momentum parameter as
@@
where $r$ is the radius of the layer in the pre-collapse core and $v_{\rm conv}$ is the convective velocity as the one-dimensional stellar evolutionary code gives according to the mixing length theory (MLT). Therefore, the specific angular momentum of the gas formed by several cells will be smaller than $j$, the value of one convective cell. If there is sufficient amplification of the specific angular momentum fluctuations by the spiral-SASI or another type of instability, then, at sporadic periods, the specific angular momentum of the accreted gas is sufficiently large to form an accretion disk (or an accretion belt) around the NS. Therefore, for launching jittering jets, the JJEM conjectures that the value of $j$ can be smaller than $j_{\rm d,NS}$. This conjecture has not been shown yet in independent simulations, but 3D simulations of the delayed-neutrino mechanism have hinted at a factor $\simeq 3$ scaling between one-dimensional MLT convection and that in hydrodynamical simulations (e.g., Fields & Couch 2020), further softening this assumption. We follow Shishkin & Soker (2022) who study FeCCSNe, and assume the constraint for ECSNe to be the same, i.e., $j\gtrsim 0.1-0.25j_{\rm d,NS}\simeq 2\times 10^{15}-5\times 10^{15}{~{}\rm cm}^{2}{~{}\rm s}^{-1}$. Earlier studies, in particular Shishkin & Soker (2022), concluded that progenitors of FeCCSNe have sufficiently large angular momentum fluctuations, i.e., $j>2\times 10^{15}{~{}\rm cm}^{2}{~{}\rm s}^{-1}$, in their convective zones near silicon burning and/or oxygen burning to support the JJEM.
To summarize, we emphasize that the criterion of $j\gtrsim 0.1-0.25j_{\rm d,NS}\simeq 2\times 10^{15}-5\times 10^{15}{~{}\rm cm}^{2}{~{}\rm s}^{-1}$ is at present an assumption of the JJEM. We here follow this earlier assumption that has been applied only to FeCCSNe (e.g., Shishkin & Soker 2021 based on Papish et al. 2015), and apply it to ECSNe. It will have to be confirmed by future studies. On the other hand, the JJEM has received very strong support from studies in the last year that identified point-symmetric morphologies in several CCSNe (e.g., Soker 2024b, a), in particular in the iconic CCSN remnant Cassiopeia A (Bear & Soker 2024). The comparison of CCSN remnants with jet-inflated structures in clusters and groups of galaxies strongly suggests that pairs of jets also shaped many CCSN remnants, and in several of these by two and more pairs of opposite jets (Soker 2024a). As stated above, these types of morphologies are expected in the JJEM and pose severe challenges to the neutrino-explosion mechanism. Nonetheless, the formation of accretions disks from fluctuations that add up to a total zero angular momentum is the main theoretical challenge of the JJEM. While previously quantified (Gilkis & Soker, 2014, 2015), this reduction in the stochastic angular momentum of the material reaching the spiral SASI modes remains a hindering factor. Confronting this challenge requires future studies to use magneto-hydrodynamical simulations constructed specifically to search for the formation of intermittent accretion disks from turbulent convective matter.
@@ 3 Numerical Scheme
We use a custom large nuclear network of 132 isotopes up to 60Zn (see Fig. 7 in Appendix) to account for weak reactions, namely, electron capture and Urca processes, that arise after the formation of the ONeMg core (e.g., Jones et al. 2013; Takahashi et al. 2013). The weak rates for the atomic mass of A=20 to A=28 isotopes tabulated by Suzuki et al. (2016) are used to account for reactions relevant to high-density ONeMg cores.
@@
Inlists and a selected profile of the M88 (Section 4) model are available online, at https://doi.org/10.5281/zenodo.10827141 (catalog doi:10.5281/zenodo.10827141).
@@ 4.1 General properties
We examine the properties of the convection layer in the outer parts of the core and above the core as close as we can get the simulation to electron-capture core collapse. We do not study the entire chain of events that lead to the collapse of the core which can be found in other studies (e.g., Jones et al. 2013; Woosley & Heger 2015; Takahashi et al. 2013; Nomoto 1987; Miyaji et al. 1980; Leung & Nomoto 2019; Zha et al. 2019; Limongi et al. 2024). These, and other studies, find that in the late evolutionary stages of an ECSN progenitor close to collapse, processes driven by the weak nuclear force, such as the Urca processes and electron capture (mainly onto 24Mg), remove energy from the center of the core. As a result, the core contracts, and oxygen and neon start to burn off-center, producing heavier elements, mainly 28Si and 32S. The O+Ne off-center burning can be described as a laminar flame propagating inwards (for treatments of laminar flames see, e.g., Timmes & Swesty 2000; Leung & Nomoto 2019; Jones et al. 2013, 2016). As the deflagration flame approaches the center, it triggers further electron capture reactions (mainly onto 20Ne) and induces collapse (Takahashi et al. 2013; Woosley & Heger 2015). The burning can reach iron peak elements (e.g., Leung & Nomoto 2019; Takahashi et al. 2013), but ECSNe explode before they develop a massive iron core. Note that during this process, the contraction of the core is in constant competition with oxygen or sometimes neon deflagration, which could lead to a thermonuclear runaway (e.g., Miyaji et al. 1980; Nomoto 1987; Leung & Nomoto 2019).
@@
Due to numerical difficulties, we do not reach actual core collapse. Based on the core composition at the last simulation time, and comparison to Doherty et al. (2015), Jones et al. (2013), Stockinger et al. (2020), and Limongi et al. (2024) that we referred to above, we infer that the relevant convective zones (see below) above the core in our last stellar model are not likely to change much until the full collapse. There is still some evolution of the inner core to collapse, like towards nuclear statistical equilibrium (NSE) in the inner core. The $8.8~{}M_{\odot}$ model with a solar composition that Stockinger et al. (2020) present just at core collapse has an inner mass of $0.45~{}M_{\odot}$ in NSE, and a CO core radius of only $\simeq 1200{~{}\rm km}$. Also, they do not have an extended helium-rich layer as we have here, of about $1~{}M_{\odot}$. However, the exact structure depends on the helium core mass. Hillebrandt et al. (1984) find that in their simulation starting with a helium core mass of $2.2~{}M_{\odot}$, at a time of 100 years before the collapse, there is a very thin helium-rich layer, while when they start with a helium core mass of $2.4~{}M_{\odot}$ there is a thick helium layer, $\simeq 1~{}M_{\odot}$ at that time. Our helium core mass is $\simeq 2.4~{}M_{\odot}$, and therefore we expect to have a thick helium layer, as we do get here.